### v1

# Auriga Streams III: the mass-metallicity relation does not rule out tidal mass loss in Local Group satellites

Abstract
Context. The mass-metallicity relation is a fundamental galaxy scaling law that has been extended to the faintest systems in the Local Group. The small scatter in this relation has been used to argue against severe tidal mass loss in Local Group satellites.
Aims. We aim to show that the observed scatter is consistent with tidal mass loss seen in cosmological hydrodynamical zoom-in simulations, using simulated Milky Way-mass haloes from the Auriga project as an example.
Methods. We analyze the stellar masses and mean metallicites for all accreted systems in Auriga. We compute these quantities two ways: considering the total system (bound + lost material) and only considering the bound progenitor. For both cases we fit the mass-metallicity relation and consider properties that drive scatter about these relations.
Results. Accreted systems in Auriga have a tight relation between total stellar mass and metallicity, with scatter at a fixed stellar mass driven by age. When only considering the bound progenitor, the tidally evolved mass-metallicity relation has similar scatter ($\sim$0.27 dex) as observed for the Milky Way and M31 satellites ($\sim$0.23 dex). Satellites that lie above the relation have experienced substantial mass loss and typically have low metallicity for their total stellar mass. Even satellites that fall exactly on the evolved relation can lose over half of their stellar mass. Only satellites that are substantially below the evolved relation are reliably intact.
Conclusions. The small scatter in the observed mass-metallicity relation is compatible with tidal mass loss. Based on their offset from the observed relation, we predict which Milky Way and M31 satellites have experienced substantial mass loss. These predictions will be tested with deep resolved-star observations from upcoming facilities including Rubin LSST, Euclid, and Roman.
Key Words.:
Local Group – Galaxy: halo – Galaxies: formation – Galaxies: evolution
1 Introduction
The mass-metallicity relation is a fundamental scaling law connecting the stellar mass of a galaxy to its metal contents (e.g. Tremonti et al., 2004; Kewley & Ellison, 2008; Mannucci et al., 2010; Maiolino & Mannucci, 2019). The relation is governed by the complex interplay between star formation, stellar evolution, accretion, and baryonic feedback. It has been extended to the faintest galaxies within the Local Volume (Simon & Geha, 2007; Kirby et al., 2008, 2013, 2020), many of which are satellites of the Milky Way and M31.
The observed scatter [^1] We adopt the RMSE, $\sigma_{\text{scatter}}\equiv\left[\sum\left(\text{[Fe/H]}_{\text{obs}}-\text{[Fe/H]}_{\text{model}}{}\right)^{2}/N\right]^{1/2}$. in this relation is remarkably small, $\sim$0.15 to 0.25 dex depending on adopted measurements. It is conventionally argued (e.g. Simon & Geha, 2007; Simon, 2019) that this small scatter indicates that the Local Group satellites have not experienced substantial mass loss. This is because individual systems would move toward smaller mass and higher mean metallicity if the system begins with a negative radial metallicity gradient and low-metallicity stars in the outskirts are preferentially lost. This would tend to increase scatter in the correlation.
Many cosmological hydrodynamical zoom-in simulations of Milky Way-mass galaxies claim to match the mass-metallicity relation (Genina et al., 2019; Grand et al., 2021; Applebaum et al., 2021), often with similar or smaller scatter than the observations. Recent analyses of such simulations (Shipp et al., 2023, 2025; Riley et al., 2025) show that their satellites have experienced much more tidal disruption than previously understood. How do these simulations produce mass-metallicity relations with small scatter, despite such a high level of mass loss? Is the observed mass-metallicity relation indicative of an intrinsic relation, or possibly one that has evolved due to tides?
Our goal in this work is to examine these questions using the Auriga simulations. We restrict our analysis to satellites that reach $M_{\star}\gtrsim 5\times 10^{5}$ M, excluding the ultrafaint systems that may not follow the same mass-metallicity relation (Fu et al., 2023) due to higher sensitivity to supernova feedback strength and other galaxy formation physics (Agertz et al., 2020).
2 Disrupting satellites in Auriga
We use the suite of Milky Way-mass haloes from the Auriga project (Grand et al., 2017, 2024). These haloes were selected from the EAGLE dark-matter-only 100 Mpc box (Schaye et al., 2015) to have $M_{\text{200c}}=1-2\times 10^{12}$ M and satisfy an isolation criterion. They were resimulated with the moving-mesh code Arepo (Springel, 2010; Pakmor et al., 2016) and a galaxy formation model detailed in Grand et al. (2017, 2024). The Auriga model produces spiral disc galaxies that are broadly consistent with a number of observations including stellar masses, sizes, and rotation curves (Grand et al., 2017), H i gas distributions (Marinacci et al., 2017), stellar disc warps (Gómez et al., 2017), stellar bars (Fragkoudi et al., 2020, 2025), satellite galaxies (Simpson et al., 2018), stellar haloes (Monachesi et al., 2019), and magnetic fields (Pakmor et al., 2017). The simulations are described in further detail in Grand et al. (2017, 2024).
[…]
The Auriga simulations trace individual abundances for nine elements: H, He, C, O, N, Ne, Mg, Si, and Fe. The abundances of individual star particles are inherited from the gas cell they are born from. We define the metallicity of a star particle as
$\text{[Fe/H]}=\log_{10}\left(\frac{M_{\text{Fe}}}{M_{\text{H}}}\right)-\log_{10}\left(\frac{55.845}{1.008}\right)-(7.46-12),$ (1)
where $M_{X}$ is the ratio of mass in species $X$ to the total mass of the parent gas cell, the second term converts mass to number density using the atomic mass of each species, and the third term sets $\text{[Fe/H]}_{\odot}\equiv 0$ (Asplund et al., 2021). When computing the mean metallicities for an accretion event, we follow conventions for resolved star observations in the Local Group. In particular, we weight each star particle by its present-day stellar mass and compute the mean of [Fe/H] as defined in Equation 1 (see Section 3.1 of Escala et al., 2018). For each system [^2] We remove 13 systems (of 419 total) that are phase-mixed with intrinsic $M_{\star}\lesssim 10^{7}$ M and [Fe/H] $\gtrsim$ -0.9. These are contaminated by merger-driven starbursts, becoming outliers in mass-metallicity-age. we compute the mean metallicity two ways: across all stars formed in that system (‘total’) and only those bound to the progenitor (‘bound’). We refer to the resulting mass-metallicity relations as ‘intrinsic’ and ‘(tidally) evolved’, respectively. We denote the fraction of total stellar mass that remains bound to the progenitor as $f_{\text{bound}}$ [^3] We refer to Sections 1 and 3 of Paper I for adopted nomenclature and definitions (e.g., ‘bound’, ‘progenitor’, ‘total stellar mass’ of a system)..
We caution that the Auriga model produces central galaxies, satellites, and stellar haloes that are more metal rich than in observations (Monachesi et al., 2019; Grand et al., 2021; Kizhuprakkat et al., 2024). We focus on the evolution of the mass-metallicity relation as Auriga satellites lose stellar mass, which depends on relative changes. In Section 4 we compare the offsets from the mass-metallicity relation for both Auriga and observations, but only after normalizing by the scatter about their respective relations, which also mitigates this issue.
3 Intrinsic vs. evolved mass-metallicity relation
Figure 1: The intrinsic mass-metallicity relation for accreted systems in Auriga. Squares indicate systems that have been fully disrupted ($f_{\text{bound}}=0$) while crosses indicate systems that still have a bound progenitor at the present day ($f_{\text{bound}}>0$). The scatter at fixed total stellar mass is driven by age (colour of points), parametrised here as $t_{90}$.
In Figure 1 we present the intrinsic mass-metallicity relation for all accreted systems in Auriga, regardless of whether or not they have a bound progenitor at the present day. The scatter at fixed total stellar mass is driven by age [^4] Parametrised as the lookback time by which 90 per cent of stellar mass formed and denoted as $t_{90}$.; older objects at a given stellar mass have lower metallicity and are more likely to be fully disrupted. This naturally arises from a mass-metallicity relation that evolves with redshift due to self enrichment (e.g. Ma et al., 2016; Torrey et al., 2019), combined with environmental quenching and tidal disruption experienced by objects from a cosmological accretion history. We also find that fully disrupted systems have higher $\alpha$/Fe] than those that survive to the present day, a familiar result in both simulations (Robertson et al., 2005; Font et al., 2006; Fattahi et al., 2020; Grimozzi et al., 2024; Pathak et al., 2025) and observations (Ji et al., 2020; Hasselquist et al., 2021; Naidu et al., 2022) that is driven by the same physical processes.
Figure 2: The intrinsic and tidally evolved mass-metallicity relations for Auriga satellites that have $f_{\text{bound}}>0$. Individual systems are shown both considering bound progenitors (circles) and the total system (crosses). We provide linear fits to both bound (solid red) and total (dashed black) cases, along with the vertical scatter about these relations. We only include systems that have 10 or more bound star particles at the present day.
In Figure 2 we present both the intrinsic and evolved mass-metallicity relations for surviving Auriga satellites (i.e. $f_{\text{bound}}>0$). Every satellite is represented in the Figure twice: crosses indicate values when considering bound and lost material (at the same locations as in Figure 1), while the circles only consider stars that remain bound to the progenitor at the present day. As satellites disrupt, they move towards lower (bound) stellar mass and higher mean metallicity. The increase in mean metallicity is a result of negative radial metallicity gradients in Auriga [^5] We discuss ignoring metallicity gradients in Appendix A. (Khoperskov et al., 2023; Orkney et al., 2023; Carrillo et al., 2024) – as lower metallicity stars in the outer regions are removed first, the remaining bound progenitor has a higher mean metallicity than the overall system. We note that negative metallicity gradients are common in observed low-mass galaxies in the Local Group and beyond (Kirby et al., 2011; Leaman et al., 2013; Taibi et al., 2022; Barbosa et al., 2025).
We fit linear models to both the intrinsic and tidally evolved data for systems that have $f_{\text{bound}}>0$. Due to tidal mass loss and negative metallicity gradients, the evolved relation sits higher and has a larger scatter (0.27) than the intrinsic relation (0.14). The tidally evolved relation, which only considers the bound progenitor, is the simulated analogue to the measurements reported in Kirby et al. (2013) and similar analyses.
The scatter in the tidally evolved relation is primarily driven by mass loss. At fixed bound stellar mass, satellites with high $f_{\text{bound}}$ are generally below the relation, while satellites with low $f_{\text{bound}}$ are above the relation. Even satellites that fall exactly on the relation can lose substantial amounts of stellar mass. We also note that individual points ‘invert’ their relationship to the mass-metallicity relation. Satellites that have low $f_{\text{bound}}$ move in the mass-metallicity plane from being below the intrinsic relation to being above the tidally evolved relation. Satellites that have high $f_{\text{bound}}$ begin above the intrinsic relation and remain in the same location of the mass-metallicity plane, but the relation shifts such that they end up below the tidally evolved relation.
Finally, we note that there is a collection of satellites with $f_{\text{bound}}\simeq 1$ that fall on or slightly above the tidally evolved relation. These systems are star forming, recently accreted satellites that have the highest possible metallicity for their total stellar mass (see Figure 1). Such systems are not likely to be considered simultaneously with satellites that show no evidence of recent star formation (e.g. Kirby et al., 2013). We show them in Figure 2 but exclude them elsewhere in this work.
4 The Local Group satellites
In Section 3 we established a connection between how disrupted a satellite in Auriga is and where it sits in the mass-metallicity plane relative to the tidally evolved relation. Now we connect this result to the observed satellites of the Milky Way and M31. We consider satellites that are brighter than $M_{V}=-7.7$ (Simon, 2019) and have literature [Fe/H] measurements derived from spectroscopy. We adopt values reported in Pace (2024) for all stellar mass measurements [^6] In practice, Pace (2024) take literature $V$-band luminosities and adopt a uniform $M_{\star}/L_{V}=2$. and augment their spectroscopic measurements with values for high-mass satellites from either McConnachie (2012) or Kirby et al. (2013). Specific measurements and their original sources are listed in Appendix B.
Figure 3: The mass-metallicity relation for observed satellites of the Milky Way (blue points) and M31 (orange points). We provide a linear fit to these satellites (solid) and the relation from Kirby et al. (2013) (dashed). We also highlight individual satellites known or believed to be disrupting and their reconstructed ‘total’ values (star symbols, see text for details).
In Figure 3 we present these measurements along with the relation reported in Kirby et al. (2013) [^7] For consistency with our adopted stellar masses, we use the luminosity-metallicity relation reported in Equation 3 of Kirby et al. (2013) and apply the same $M_{\star}/L_{V}=2$.. We note that the fit reported in Kirby et al. (2013) includes ultrafaint systems down to $10^{3}$ M, which do not appear to continue the relation and instead flatten at [Fe/H] $\sim$ 2.6-2.6 (Fu et al., 2023). For a consistent treatment, we fit a new relation as done for the Auriga satellites (Figure 2). This fit reduces the scatter from 0.24 to 0.23, which is comparable to the 0.27 for bound satellites in Auriga.
Figure 4: Top panel: $f_{\text{bound}}$ vs. scaled offset from the evolved mass-metallicity relation for Auriga satellites. The dashed lines are the boundaries between the five broad categories of mass loss detailed in Section 4 and Appendix B. Bottom panel: histogram of scaled offsets for both Auriga and Milky Way and M31 satellites from their respective mass-metallicity relations.
We compute the vertical (i.e. at fixed mass) offset from the fit $\Delta$[Fe/H] for each satellite, which is positive (negative) for satellites that sit above (below) the relation. In the top panel of Figure 4 we present these offsets and show they broadly correlate with $f_{\text{bound}}$ for the Auriga satellites, as seen in Figure 2.
We can leverage this connection between $f_{\text{bound}}$ and $\Delta$[Fe/H] to predict which Local Group satellites are actively disrupting. This is detailed in Appendix B, but in brief, we assign a Class 1 through 5 based on the offset $\Delta$[Fe/H] normalised by the $\sigma_{\text{scatter}}$ about the relation. The boundaries between the different classes are shown as dashed lines in Figure 4 and largely capture different categories of mass loss in Auriga, from unlikely to be losing mass (Class 1) to guaranteed to have experienced extreme disruption (Class 5). In particular, we highlight the following quiescent systems predicted to be disrupting at high confidence:
  • Class 3 (0.05 ¡ $f_{\text{bound}}$ ¡ 0.8): Fornax, Leo II, NGC 147, And I, II, III, VI, XIX, XXI, XXVIII, XXIX
  • Class 4 ($f_{\text{bound}}<0.2$): And XVII, And XVIII, Cassiopeia II
  • Class 5 ($f_{\text{bound}}<0.05$): Sagittarius, And XV
The offset $\Delta$[Fe/H] and predicted disruption Class for individual Local Group satellites are reported in Appendix B.
It may appear surprising that so many Milky Way satellites could be significantly disrupted, especially given current detections. However, the tidal tails emanating from these systems are likely too faint to be detected in current resolved-star imaging, so the satellites would appear intact (Shipp et al., 2023). These low surface brightness features may be visible by upcoming facilities including Rubin LSST, Euclid, and Roman.
To illustrate how far satellites can move in the mass-metallicity plane, in Figure 3 (star symbols) we provide reconstructions of the ‘total’ system for three prominent examples of disrupting Milky Way satellites: Sagittarius, Antlia II, and Crater II. For Sagittarius we adopt an $f_{\text{bound}}$ == 0.19 (Niederste-Ostholt et al., 2010) and assume that the progenitor has mean [Fe/H] == 0.58-0.58, while the tails are more metal poor with [Fe/H] == 1.07-1.07 (Limberg et al., 2022; Cunningham et al., 2024). For Antlia II and Crater II we adopt modeling estimates of $f_{\text{bound}}$ == 0.1 (Sameie et al., 2020) and 0.3 (Sanders et al., 2018) respectively [^8] Ji et al. (2021) argue against substantial mass loss for Antlia II and Crater II based on offset from the Kirby et al. (2013) relation. We assign both objects to Class 2 where $f_{\text{bound}}\gtrsim 0.2$ is possible, such that the Sameie et al. (2020) Antlia II estimate of 0.1 is difficult to match but the Sanders et al. (2018) Crater II estimate of 0.3 is feasible. and assume there is no shift in metallicity given the lack of a detected metallicity gradient (Ji et al., 2021).
Even though these estimates are simple, they illustrate (based on independent modeling) how far individual satellites can travel in the mass-metallicity plane. In addition, it is clear that their ‘total’ mass and metallicity should not be compared to the mass-metallicity relation of satellites, but to the (unknown) intrinsic relation (as highlighted in Figure 2). It is also intriguing that these reconstructions move their respective satellites near the lowest edge of the satellite mass-metallicity relation, suggesting that this may be near where the intrinsic relation lies.
Finally, when estimating the stellar mass of streams a common technique (e.g. Li et al., 2022) is to measure the stream’s metallicity and then adopt a mass-metallicity relation, often from Kirby et al. (2013). We emulate this method for Auriga systems by computing the mean metallicity of unbound stars and converting this into a stellar mass using the evolved relation (solid line in Figure 2). We find this underestimates the true total mass with $\log_{10}(M_{\ast}^{\text{est}}/M_{\ast}^{\text{true}})=-0.34^{+0.38}_{-0.59}$ for systems with a progenitor at the present day and $\log_{10}(M_{\ast}^{\text{est}}/M_{\ast}^{\text{true}})=-0.57^{+0.33}_{-0.34}$ for systems with no progenitor. The latter is impacted by redshift evolution of the intrinsic relation (Figure 1). While this technique is an improvement over methods using star counts for only the chunk of stream that is detected, we recommend such estimates should be considered lower bounds on the total mass of the system.
5 Concluding remarks
Accreted systems in Auriga follow established correlations between stellar mass, metallicity, and star formation history (Figure 1) that, upon accretion and disruption, imprint a tidally evolved mass-metallicity relation that differs from the intrinsic one (Figure 2). The scatter about this tidally evolved relation is similar to that observed for Local Group satellites (Figure 3) and offset from the relation predicts which satellites have experienced substantial mass loss (Figure 4). The mass-metallicity relation for Local Group satellites is not only compatible with tidal mass loss, but provides further evidence that a wealth of faint streams awaits discovery in Rubin, Euclid, and Roman data.
Acknowledgements.
We thank Astha Chaturvedi, Emma Dodd, Azi Fattahi, Peter Ferguson, and Guilherme Limberg for enlightening discussions, as well as Jess Doppel and Louise Welsh for peer pressure to ‘just write the damn paper’. AHR thanks Taylor Swift’s 1989 (Taylor’s Version) and Bring Me The Horizon’s POST HUMAN: NeX GEn for providing a soundscape condusive to writing this manuscript. We have used simulations from the Auriga Project public data release (Grand et al., 2024) available on the Globus platform [^9] https://wwwmpa.mpa-garching.mpg.de/auriga/data. This research made extensive use of arXiv.org and NASA’s Astrophysics Data System for bibliographic information.
AHR was supported by a fellowship funded by the Wenner Gren Foundation, a Research Fellowship from the Royal Commission for the Exhibition of 1851, and by STFC through grant ST/T000244/1. RB is supported by the UZH Postdoc Grant, grant no. FK-23116 and the SNSF through the Ambizione Grant PZ00P2_223532. FF is supported by a UKRI Future Leaders Fellowship (grant no. MR/X033740/1). FAG acknowledges support from the ANID BASAL project FB210003, from the ANID FONDECYT Regular grants 1251493 and from the HORIZON MSCA-2021-SE-01 Research and Innovation Programme under the Marie Sklodowska-Curie grant agreement number 101086388. RJJG is supported by an STFC Ernest Rutherford Fellowship (ST/W003643/1). FM acknowledges funding from the European Union - NextGenerationEU under the HPC project ‘National Centre for HPC, Big Data and Quantum Computing’ (PNRR - M4C2 - I1.4 - CN00000013 – CUP J33C22001170001). This work was supported by collaborative visits funded by the Cosmology and Astroparticle Student and Postdoc Exchange Network (CASPEN). We acknowledge support from the DiRAC Institute in the Department of Astronomy at the University of Washington. The DiRAC Institute is supported through generous gifts from the Charles and Lisa Simonyi Fund for Arts and Sciences, Janet and Lloyd Frink, and the Washington Research Foundation.
This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
This research made use of the Python programming language, along with many community-developed or maintained software packages including: astropy (Astropy Collaboration et al., 2013, 2018, 2022), cmasher (van der Velden, 2020; van der Velden et al., 2024) cython (Behnel et al., 2011) h5py (Collette, 2013; Collette et al., 2022) jupyter (Perez & Granger, 2007; Kluyver et al., 2016) matplotlib (Hunter, 2007) numpy (Harris et al., 2020) pandas (McKinney, 2010; pandas development team, 2022) scikit-learn (Pedregosa et al., 2011; Buitinck et al., 2013; Grisel et al., 2022) scipy (Virtanen et al., 2020; Gommers et al., 2022). We also thank the maintainers of the arepo-snap-util package. Parts of the results in this work make use of the colormaps in the cmasher package. Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2024).
For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Appendix A Ignoring metallicity gradients
Figure 5: Similar to Figure 2 but excluding the effect of radial metallicity gradients.
Many of the key results in this analysis for Auriga are driven simultaneously by both tidal mass loss and by initial negative radial metallicity gradients. Due to these gradients, low metallicity stars are preferentially lost from the outskirts, raising the mean metallicity of the bound progenitor. While we prefer to present results that include both effects, it can be interesting to see if our results qualitatively hold when ignoring the effect of radial gradients.
In Figure 5 we show a version of Figure 2 that ignores the metallicity evolution of the bound progenitor by adopting the metallicity of the total system. In other words, individual points move exclusively to lower stellar mass. We still find that the mean relation moves away from the intrinsic one, the scatter about the relation increases, and that individual points ‘invert’ their relationship to the relation.
Appendix B Disruption prediction for observed satellites
Class Definition $f_{\text{bound}}^{\text{Au}}$ $F^{\text{Au}}_{\text{intact}}$ $F^{\text{Au}}_{\text{<0.8}}$ $N^{\text{obs}}_{\text{sat}}$
$\Delta\text{[Fe/H]}/\sigma\in$
1 $(-\infty,-1]$ $0.92^{+0.07}_{-0.46}$ 0.38 0.31 5
2 (1,0](-1,0] $0.81^{+0.18}_{-0.43}$ 0.21 0.47 17
3 $(0,1]$ $39^{+0.42}_{-0.26} 0.0$7 0.83 12
4 $(1,2]$ $06^{+0.05}_{-0.02} 0.0$0 1.00 3
5 $(2,\infty)$ $00^{+0.01}_{-0.00} 0.0$0 1.00 2
[…]
In Section 4 we introduce a broad classification for how disrupted Auriga satellites are, based on their offset from the tidally evolved mass-metallicity relation $\Delta$[Fe/H] normalised by the scatter about the relation $\sigma_{\text{scatter}}$. The exact boundaries are listed in Table 1 and separate into the following qualitative descriptions:
  • Class 1: unlikely to have experienced any mass loss
  • Class 2: mass loss down to $f_{\text{bound}}\sim 0.2$ is possible but disfavoured
  • Class 3: substantial mass loss is favoured, with most satellites having $0.05<f_{\text{bound}}<0.8$
  • Class 4: no intact satellites, $f_{\text{bound}}<0.2$ strongly preferred
  • Class 5: no intact satellites, only $f_{\text{bound}}<0.05$
Table 1 also lists quantitative metrics for how disrupted Auriga satellites are in each Class. As shown in Figure 4, both the highest level of disruption that is possible and the fraction of satellites that are substantially disrupted increase with ascending Class order. We caution that these predictions are less robust for systems that have recent star formation (see final paragraph of Section 3).
Name $\log_{10}(M_{\star}/\text{M${}_{\odot}$})$ [Fe/H] $\Delta$[Fe/H] Class References
Milky Way satellites
Antlia II 6.13 -1.9 -0.15 2 [1,1]
Canes Venatici I 5.73 -1.91 -0.02 2 [2,3]
Carina 6.01 -1.8 -0.0 2 [2,3]
Crater II 5.51 -2.16 -0.19 2 [4,1]
Draco 5.78 -2.0 -0.13 2 [2,3]
Fornax 7.59 -1.07 0.18 3 [2,3]
Leo I 6.96 -1.48 -0.02 2 [2,3]
Leo II 6.12 -1.68 0.08 3 [2,3]
LMC 9.46 -0.5 0.1 3 [5,6]
Sagittarius 7.63 -0.53 0.7 5 [7,3]
Sculptor 6.54 -1.73 -0.12 2 [2,3]
Sextans 5.72 -1.97 -0.08 2 [2,3]
SMC 8.95 -1.0 -0.22 2 [5,8]
Ursa Minor 5.78 -2.13 -0.26 1 [2,9]
M31 satellites
Andromeda I 6.77 -1.51 0.02 3 [10,11]
Andromeda II 6.92 -1.47 0.01 3 [10,12]
Andromeda III 6.03 -1.75 0.04 3 [10,11]
Andromeda V 5.95 -1.84 -0.03 2 [10,11]
Andromeda VI 6.75 -1.5 0.04 3 [13,14]
Andromeda VII 7.35 -1.37 -0.04 2 [13,11]
Andromeda IX 5.69 -2.03 -0.12 2 [10,12]
Andromeda XIV 5.69 -2.23 -0.33 1 [10,12]
Andromeda XV 5.58 -1.43 0.51 5 [10,12]
Andromeda XVII 5.35 -1.7 0.32 4 [10,14]
Andromeda XVIII 5.9 -1.53 0.3 4 [10,15]
Andromeda XIX 6.25 -1.5 0.21 3 [10,16]
Andromeda XXI 5.81 -1.7 0.16 3 [10,17]
Andromeda XXIII 6.14 -2.2 -0.45 1 [10,14]
Andromeda XXV 5.87 -1.9 -0.06 2 [10,18]
Andromeda XXVII 5.39 -2.1 -0.09 2 [19,14]
Andromeda XXVIII 5.64 -1.84 0.08 3 [20,20]
Andromeda XXIX 5.5 -1.9 0.07 3 [20,20]
Cassiopeia II 5.33 -1.7 0.33 4 [10,14]
Cassiopeia III 7.19 -1.7 -0.32 1 [21,22]
Lacerta I 6.72 -2.0 -0.45 1 [21,22]
NGC 147 8.17 -0.83 0.22 3 [7,23]
NGC 185 8.18 -1.12 -0.08 2 [7,23]
NGC 205 8.84 -0.92 -0.1 2 [7,23]
Perseus I 5.78 -2.0 -0.13 2 [21,22]
Table 2: Observational data (compiled by Pace 2024) presented in Figure 3 along with the vertical offset from the mass-metallicity relation ($\Delta$[Fe/H]) and a Class indicating the potential level of (tidal) disruption experienced by the system (see text in Appendix B for details). We also list references for each measurement, which correspond to the following citations: (1) Ji et al. (2021) (2) Muñoz et al. (2018) (3) Simon (2019) (4) Torrealba et al. (2016) (5) de Vaucouleurs et al. (1991) (6) Carrera et al. (2008) (7) McConnachie (2012) (8) Parisi et al. (2010) (9) Pace et al. (2020) (10) Martin et al. (2016) (11) Kirby et al. (2020) (12) Wojno et al. (2020) (13) McConnachie & Irwin (2006) (14) Collins et al. (2013) (15) Kvasova et al. (2024) (16) Cullinane et al. (2024) (17) Collins et al. (2021) (18) Charles et al. (2023) (19) Richardson et al. (2011) (20) Slater et al. (2015) (21) Rhode et al. (2023) (22) Martin et al. (2014) (23) Kirby et al. (2013).
We then apply the same exercise to the observed Local Group satellites to predict how much tidal disruption they have experienced. We begin with compilations of their stellar masses and mean metallicities (Figure 3), then fit a relation in the same manner as done for the Auriga satellites. Our fit takes the following form:
$\text{[Fe/H]}=-1.80+0.35\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{\leavevmode\nobreak\ M${}_{\odot}$}}\right)$ (2)
The equivalent for Kirby et al. (2013), taking the luminosity-metallicity relation reported in their Equation 3 and assuming $M/L_{V}=2$, is
$\text{[Fe/H]}=-1.77+0.29\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{\leavevmode\nobreak\ M${}_{\odot}$}}\right)$ (3)
We compute each satellite’s offset from the relation (Equation 2) normalized by the reported scatter ($\sigma_{\text{scatter}}=0.23$). These are then directly compared to the Class boundary definitions in Table 1. The stellar masses, metallicities, offsets from the relation in Equation 2, and resulting Class assignments are indicated in Table 2.
We note that there are three M31 satellites (M 32, IC 10, and LGS 3) brighter than $M_{V}$ = 7.7-7.7 omitted from our analysis because they lack literature spectroscopic metallicities in the compilations we consider. The available isochrone or RGB colour metallicities compiled by McConnachie (2012) suggest that IC 10 and LGS 3 are not candidates for heavy disruption under our framework (Class 1 or 2 depending on the adopted relation). However, the available data for M 32 gives it a very high metallicity ($\text{[Fe/H]}=-0.25$; Grillmair et al. 1996) for its stellar mass. This would place M 32 in Class 5, comparable to Sagittarius.